DIRC Operations Report
Wednesday 2007/6/6 to Wednesday 2007/7/11
Oncall: Alejandro Perez, Bradley Wogsland, Justine Serrano
Fast Monitoring: Alejandro Perez, Bradley Wogsland, Justine Serrano, Jose Benitez
Luminosities this week
PEP peak lumi: ~1.02 x 1034
(last week: ~1 x 1034)
PEP delivery: 3.92 fb-1
(last weeks: 3.29 fb-1)
BaBar integrated lumi:3.8 fb-1
(last weeks: 3.17 fb-1)
Run 6 total delivery: 58.29 fb-1
Run 6 total BaBar recorded: 56.02 fb-1 total, including 4.18 fb-1 off-peak data
PEP activities/issues since last DIRC meeting
06/11: Switched to Offpeak running.
Hot weather has disrupted operations, limited data integration.
06/15: Magnet Power Supply tripped around 6.30am. Recoverred by 8am.
06/19: After taking 4.18 fb-1 of offpeak data PEP switched to Onpeak running.
06/21 to 06/22: ROD. They replace a bellow in arc3 which was producing a lot of HER aborts.
06/27: ROD. PEP people remove two flexible flanges in arc5. Beams back in the night.
07/09: Vacuum problem since several days, finally a vacuum leak was found near backward Q5.
BaBar activities/issues since last DIRC meeting
06/11: UPS battery replacement.
During the week, some subsystems have raised temperature thresholds because of hot weather.
06/15: HVAC rebooted their control computer and all the central chiller went off (accidentally).
06/21: ROD. Tried to fix a EH cooling problem but the spare didn't fit.
06/27: ROD. Two FEAs from the DCH has been changed succesfully.
DIRC activities/issues since last DIRC meeting
06/07: New problem with 8 flaky channels in sector 3.
06/11: Masked 8 noisy channels (7 in sector 6 and 1 in sector 8)
06/14: A flaky temperature sensors (in sector 10) started to give many yellow alarms due to heat wave. We raised the inlet temperature yellow alarm threshold for this sector (from 34 to 35 degrees).
06/15: Due to central chiller problem the DIRC chiller tripped and FEE went off. DIRC recovered fine after HVAC people turned central chillers back on.
06/19: 8 noisy channels masked out (DFB 0, sector 0).
06/21: ROD activities:
06/22: After the ROD, HV group 1 in sector 0 tripped because of the Xmas tree. We set the runnable value at 900V so that it didn't trip anymore.
06/27: ROD Activities:
- We turned off the HV, FEE crates and the chiller for EH work.
- Mike replaced a connector card at the back of the chiller power siam.
- While BaBar was starting to take cosmics, we lost the monitoring of sector 5.
Bad background conditions since the 3rd of July which is due to a vacuum leak. This causes a lot of red alarms in the scalers and DIRC dead time. We raised the alarm thresholds :
- X-Trees in sector 0 and sector 7 unplugged.
- Replaced DFB #1 in sector 6.
- Replaced DCC in sector 5.
- Before the ROD, one calibration failed. This was due to a bad board in the master timing crate, Chris O'Grady replaced it in the afternoon.
- scalers: yellow @ 600kHz, red @ 700 kHz (previously: yellow @ 400 and red @ 600 kHz)
- rate ratios: yellow @ 3.0, red @ 4.0 (previously: yellow @ 1.5 and red @ 2.5)
- integrated scalers: yellow @ 9.0, red @ 9.5
See DIRC logbook for details
Lumi and Background this month
- Hotest scaler rates
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Replace air filters on Wienner crates.
- Investigate the crate of sector 8 which has again killed the first 8 channels of DFB 7.
| Equipment || # used || # spare (*) || Comments |
| DFB || 168 || 18
6 good (3F,4E,64,66,68,69),
3 potentially good (6E,57,9C)
9 bad (05,70,6C,78,40,88,95,87,5A) in France
| DCC || 12 || 7
3 good (12,31,5)
2 bad (25,9) in France
- Justine and Jose rotating backup & on-call commissioning duties.
- Rolf is coming on next week.